White Company has two departments, Cutting and Finishing. The company uses a job-order costing system and computes a predetermined overhead rate in each department. The Cutting Department bases its rate on machine-hours, and the Finishing Department bases its rate on direct labor-hours. At the beginning of the year, the company made the following estimates: Department Cutting Finishing Direct labor-hours Machine-hours Total fixed manufacturing overhead cost Variable manufacturing overhead per machine-hour Variable manufacturing overhead per direct labor-hour 64,000 1,400 $ 380,000 $ 444,000 8,900 55,300 $ 3.00 $ 2.75 Required 1. Compute the predetermined overhead rate for each department. 2. The job cost sheet for Job 203, which was started and completed during the year, showed the following: Department Cutting Finishing Direct labor-hours Machine-hours Direct materials Direct labor cost 87 760 $ 63 $360 $357 Using the predetermined overhead rates that you computed in requirement (1), compute the total manufacturing cost assigned to Job 203 3. Would you expect substantially different amounts of overhead cost to be assigned to some jobs if the company used a plantwide predetermined overhead rate based on direct labor-hours, rather than using departmental rates? Complete the question by entering your answers in the tabs given belovw Required 1Required 2 Required 3 Compute the predetermined overhead rate for each department. (Round your answers to 2 decimal places.) Department Cuttin Finishin Predetermined overhead rate per MH per DLH Required 2 >

Required 1 Required 2 Required 3 Would you expect substantially different amounts of overhead cost to be assigned to some jobs if the company used a plantwide predetermined overhead rate based on direct labor-hours, rather than using departmental rates? Required 2 Required3
Answer #1
1
Predetermined overhead rate:
Cutting 9.87 per MH =3+(380000/55300)
Finishing 9.69 per DLH =2.75+(444000/64000)
2
Direct materials 1120 =760+360
Direct labor cost 420 =63+357
Overhead applied 1023 =(87*9.87)+(17*9.69)
Total manufacturing cost 2563 or 2564
3
Yes, the amounts of overhead cost would be substantially different